# Kanchon Gharami, Sanjiv Kumar Sarkar, Yongxin Liu§, and Shafika Showkat Moni

### v1

@@ I Introduction
At the same time, editors face a different challenge. They want to detect when reviewers secretly use LLMs. To do this, some systems insert a small synthetic pattern inside the manuscript. A human reviewer ignores it, but an LLM often repeats or expands this pattern in the review. This creates a simple signal that reveals AI-written reports. Both attack and defence now rely on subtle interactions between the document and the LLM, which makes the problem more complex.
Although several papers explore these ideas, most prior work studies them in isolation. Some focus only on hidden-text attacks. Others analyze general weaknesses in LLM reviewers. A few propose structural detectors for hidden prompts. However, there is still no complete analysis that connects realistic attack methods, document parsing behaviour, and practical defences that work without modifying the LLM. A unified view is needed to understand how these systems behave in real reviewing environments.
This paper provides that unified view. We design a hybrid attack that combines hidden-text injection, smooth topic-shift segments, and iterative prompt refinement using a surrogate model. This reflects what a motivated attacker can realistically do. We then propose a two-layer defence. The first layer checks the structure of the document using two independent text views. The second layer checks how stable the LLM output is when the input is slightly changed. Finally, we add an editor-side trap that helps detect when reviewers use LLMs even if the manuscript is safe.
@@
  • We develop a two-layer defence that combines structural reconstruction with behavioural mutation tests, working without access to model weights or training data.
@@ III-D Security Goals
A successful defence should work against both simple and optimized attacks without requiring retraining of the LLM.
@@
Figure 2: High level overview of defence pipeline.
@@ IV Methodology
This section describes the complete attack and defence pipeline used in this work. The goal is to study how an uploaded document can influence a Large Language Model (LLM) and how the system can detect such manipulation before generating the final answer. The process begins with a document $D$, which is parsed into text $T$, and then combined with the system prompt $s$ and the user request $q$. We build a hybrid attack method inspired by recent prompt-in-content ideas and design a two-layer defence that combines structural checks with behavioural analysis. The overall flow is shown in Fig. 1 and Fig. 2.
@@ IV-A3 Iterative Prompt Optimization
IV-B Defence Model and Pipeline
The defence pipeline uses two layers. The first layer checks the structure of the document, and the second layer checks the stability of the model output. Together these steps detect both simple and adaptive attacks without modifying the LLM itself.
@@ IV-B5 Final Decision
The defence rejects a document if any structural or behavioural signal exceeds its threshold:
@@
The full methodology combines multiple attack strategies and multiple defence signals. The attack injects hidden text, uses a topic shift, and refines the instruction through iterations. The defence checks for hidden content, screens instruction-like text, tests behavioural stability, and adds an editor-side trap against LLM-generated reviews. This creates a complete and practical framework for understanding and securing AI-assisted reviewing systems.
@@ VII Conclusion
This paper examined how hidden instructions inside uploaded documents can manipulate large language models in AI-assisted peer review. We showed that the document becomes an attack surface and proposed a hybrid prompt-in-content attack that combines hidden text, topic-shift cues, and iterative refinement. To counter these threats, we introduced a two-layer defence based on structural reconstruction and behavioural stability checks, along with an editor-side trap that helps detect LLM-generated reviews.
Overall, our results show that these attacks are easy to create but require dedicated defences to detect. The proposed pipeline offers a practical path for improving the security of AI-assisted reviewing systems. Future work will explore stronger defences, more document formats, and better tools for identifying LLM-assisted reviewer behaviour.

### v2

@@ I Introduction
At the same time, editors face a different challenge. They want to detect when reviewers secretly use LLMs. To do this, some systems insert a small synthetic pattern inside the manuscript. A human reviewer ignores it, but an LLM often repeats or expands this pattern in the review. This creates a simple signal that reveals AI-written reports. Both attack and defense now rely on subtle interactions between the document and the LLM, which makes the problem more complex.
Although several papers explore these ideas, most prior work studies them in isolation. Some focus only on hidden-text attacks. Others analyze general weaknesses in LLM reviewers. A few propose structural detectors for hidden prompts. However, there is still no complete analysis that connects realistic attack methods, document parsing behavior, and practical defenses that work without modifying the LLM. A unified view is needed to understand how these systems behave in real reviewing environments.
This paper provides that unified view. We design a hybrid attack that combines hidden-text injection, smooth topic-shift segments, and iterative prompt refinement using a surrogate model. This reflects what a motivated attacker can realistically do. We then propose a two-layer defense. The first layer checks the structure of the document using two independent text views. The second layer checks how stable the LLM output is when the input is slightly changed. Finally, we add an editor-side trap that helps detect when reviewers use LLMs even if the manuscript is safe.
@@
  • We develop a two-layer defense that combines structural reconstruction with behavioural mutation tests, working without access to model weights or training data.
@@ III-D Security Goals
A successful defense should work against both simple and optimized attacks without requiring retraining of the LLM.
@@
Figure 2: High level overview of defense pipeline.
@@ IV Methodology
This section describes the complete attack and defense pipeline used in this work. The goal is to study how an uploaded document can influence a Large Language Model (LLM) and how the system can detect such manipulation before generating the final answer. The process begins with a document $D$, which is parsed into text $T$, and then combined with the system prompt $s$ and the user request $q$. We build a hybrid attack method inspired by recent prompt-in-content ideas and design a two-layer defense that combines structural checks with behavioural analysis. The overall flow is shown in Fig. 1 and Fig. 2.
@@ IV-A3 Iterative Prompt Optimization
IV-B defense Model and Pipeline
The defense pipeline uses two layers. The first layer checks the structure of the document, and the second layer checks the stability of the model output. Together these steps detect both simple and adaptive attacks without modifying the LLM itself.
@@ IV-B5 Final Decision
The defense rejects a document if any structural or behavioural signal exceeds its threshold:
@@
The full methodology combines multiple attack strategies and multiple defense signals. The attack injects hidden text, uses a topic shift, and refines the instruction through iterations. The defense checks for hidden content, screens instruction-like text, tests behavioural stability, and adds an editor-side trap against LLM-generated reviews. This creates a complete and practical framework for understanding and securing AI-assisted reviewing systems.
@@ VII Conclusion
This paper examined how hidden instructions inside uploaded documents can manipulate large language models in AI-assisted peer review. We showed that the document becomes an attack surface and proposed a hybrid prompt-in-content attack that combines hidden text, topic-shift cues, and iterative refinement. To counter these threats, we introduced a two-layer defense based on structural reconstruction and behavioural stability checks, along with an editor-side trap that helps detect LLM-generated reviews.
Overall, our results show that these attacks are easy to create but require dedicated defenses to detect. The proposed pipeline offers a practical path for improving the security of AI-assisted reviewing systems. Future work will explore stronger defenses, more document formats, and better tools for identifying LLM-assisted reviewer behaviour.